the sum of two consecutive odd numbers plus the next even number is 56.write an algebraic equation that can be used to solve this problem. what is the larger odd number?
Answer by shree840(260)

x+x+2+x
+3=56
3x+5=56
3x=51
x=17
next odd number =17+2=19 ans